### v7

@@ 4.5 Comparisons With Previous Methods
It is also worth noting that although DINO achieves slightly higher performance than TinyDef-DETR on certain metrics, its training requires approximately 79.8 hours, which is significantly longer. Such excessive training time limits its practical applicability in real-world UAV-based inspection tasks, whereas TinyDef-DETR achieves a more balanced trade-off between accuracy and efficiency.
@@
4.6 Generalization Experiments for the Proposed Methods
Since small objects dominate our dataset, we further evaluate the proposed method on the VisDrone dataset (Zhu et al., 2022), a widely used UAV-based benchmark for object detection and tracking. VisDrone contains over 26,000 high-resolution images and 263 video sequences captured in diverse urban and suburban scenes, with more than 100,000 annotated objects across 10 common categories. Due to its dense object distribution, large proportion of small targets, and complex real-world conditions, VisDrone provides a challenging platform for evaluating model robustness and generalization.
@@ 4.6 Generalization Experiments for the Proposed Methods
| YOLO v10n(Wang et al., 2025) | ||||||
| YOLO v10s | ||||||
| YOLO v10m | ||||||
| YOLO v11n(Khanam and Hussain, 2024) | ||||||
| YOLO v11s | ||||||
| YOLO v11m | ||||||
| YOLO v12n(Tian et al., 2025) | ||||||
| YOLO v12s | ||||||
| YOLO v12m | ||||||
| YOLO v13n(Lei et al., 2025) | ||||||
| YOLO v13s | ||||||
@@ 5 Conclusions
Due to confidentiality agreements, part of the dataset used in this study cannot be shared publicly. However, the data presented in this study also include openly available resources from the VisDrone benchmark dataset, which can be accessed at https://github.com/VisDrone/VisDrone-Dataset, as described in Zhu et al. (2022).
@@ Acknowledgements.
- Yang et al. (2023) Yang, L.; Kong, S.; Deng, J.; Li, H.; Liu, Y. DRA-Net: A Dual-Branch Residual Attention Network for Pixelwise Power Line Detection. IEEE Trans. Instrum. Meas. 2023, 72, 1–13.
- Liu et al. (2021) Liu, X.; Miao, X.; Jiang, H.; Chen, J. Box-Point Detector: A Diagnosis Method for Insulator Faults in Power Lines Using Aerial Images and Convolutional Neural Networks. IEEE Trans. Power Delivery 2021, 36, 3765–3773.
- Yu et al. (2023) Yu, Z.; Lei, Y.; Shen, F.; Zhou, S.; Yuan, Y. Research on Identification and Detection of Transmission Line Insulator Defects Based on a Lightweight YOLOv5 Network. Remote Sensing 2023, 15, 4552.
- Zhang et al. (2025) Zhang, X.; Zhang, Y.; Shen, K.; Fu, Q.; Shen, H. FAFNet: An Overhead Transmission Line Component Detection Method Based on Feature Alignment and Fusion. IEEE Sensors J. 2025, 25, 30197–30206.
- Zhang et al. (2021) Zhang, X.; Zhang, Y.; Liu, J.; Zhang, C.; Xue, X.; Zhang, H.; Zhang, W. InsuDet: A Fault Detection Method for Insulators of Overhead Transmission Lines Using Convolutional Neural Networks. IEEE Trans. Instrum. Meas. 2021, 70, 1–12.
- Tunio et al. (2024) Tunio, N.A.; Hashmani, A.A.; Khokhar, S.; Tunio, M.A.; Faheem, M. Fault Detection and Classification in Overhead Transmission Lines through Comprehensive Feature Extraction Using Temporal Convolution Neural Network. Engineering Reports 2024, 6, e12950.
@@
- Zhu et al. (2022) Zhu, P.; Wen, L.; Du, D.; Bian, X.; Fan, H.; Hu, Q.; Ling, H. Detection and Tracking Meet Drones Challenge. IEEE Transactions on Pattern Analysis and Machine Intelligence 2022, 44, 7380–7399.

### v8

@@ 4.5 Comparisons With Previous Methods
It is also worth noting that DINO(Zhang et al., 2022)—evaluated under identical hardware and software using the official OpenMMLab Chen et al. (2019) recipe dino-4scale_r50_8xb2-12e_coco with its default 12-epoch schedule—performs worse than TinyDef-DETR across all reported metrics, while requiring 79.8 hours of training; by contrast, TinyDef-DETR, trained with our default 100-epoch schedule on the same setup, completes in under 20 hours. Although wall-clock time is environment-dependent, these figures reflect realistic, recommended settings, and TinyDef-DETR thus offers a more favorable accuracy–efficiency trade-off for real-world UAV inspection—amusingly, it seems this “longer training, better performance” effect doesn’t always make sense.
@@
4.6 Generalization Experiments for the Proposed Methods
Since small objects dominate our dataset, we further evaluate the proposed method on the VisDrone-DET 2019 benchmark, a widely used UAV dataset for object detection in images. VisDrone-DET 2019Du et al. (2019) comprises 8,599 drone-captured images with >540k annotated bounding boxes across 10 general-purpose categories (e.g., pedestrian, person, car, van, bus, truck, motor, bicycle, awning-tricycle, tricycle), split into 6,471/548/1,580 for train/val/test. Importantly, VisDrone does not include transmission-line defect classes. Therefore, our results on VisDrone (Table 5) assess cross-domain transferability rather than in-domain recognition. Owing to its dense object distribution, the high proportion of small targets, and complex real-world conditions, VisDrone provides a rigorous testbed for robustness and generalization. The strong performance we observe thus demonstrates our framework’s general efficacy for small-object detection in UAV imagery, beyond the specific domain of transmission-line inspection, which is practically valuable.
@@ 4.6 Generalization Experiments for the Proposed Methods
| YOLO 10n(Wang et al., 2025) | ||||||
| YOLO 10s | ||||||
| YOLO 10m | ||||||
| YOLO 11n(Khanam and Hussain, 2024) | ||||||
| YOLO 11s | ||||||
| YOLO 11m | ||||||
| YOLO 12n(Tian et al., 2025) | ||||||
| YOLO 12s | ||||||
| YOLO 12m | ||||||
| YOLO 13n(Lei et al., 2025) | ||||||
| YOLO 13s | ||||||
@@ 5 Conclusions
Due to confidentiality agreements, part of the dataset used in this study cannot be shared publicly. However, the data presented in this study also include openly available resources from the VisDrone benchmark dataset, which can be accessed at https://github.com/VisDrone/VisDrone-Dataset, as described in Du et al. (2019).
@@ Acknowledgements.
- Shao et al. (2025) Shao, X.; Diao, S.; Li, L.; Zhao, X.; Mei, Y.; Zhu, Z. WT-DETR: Wavelet-enhanced DETR for Robust Tiny Object Detection via Multi-Scale Feature Optimization. JOURNAL OF REAL-TIME IMAGE PROCESSING 2025, 22.
- Feng et al. (2025) Feng, D.; Xu, J.; Li, K.; Ma, Z.; Li, W.; Wang, D. FADet: A Frequency-Aware Detection Framework for Infrared Small Target Detection. IEEE JOURNAL OF SELECTED TOPICS IN APPLIED EARTH OBSERVATIONS AND REMOTE SENSING 2025, 18, 24963–24975.
- Xu et al. (2025) Xu, W.; Ding, Z.; Wang, Z.; Cui, Z.; Hu, Y.; Jiang, F. Think Locally and Act Globally: A Frequency-Spatial Fusion Network for Infrared Small Target Detection. IEEE TRANSACTIONS ON GEOSCIENCE AND REMOTE SENSING 2025, 63.
- Li et al. (2025) Li, H.; Peng, X.; Zhang, J. FSCNet: Feature Synthesis with Wavelet Coefficients for Infrared Small Target Detection. INFRARED PHYSICS & TECHNOLOGY 2025, 147.
- Wang et al. (2025) Wang, K.; Suo, B.; Zhang, G.; Yang, Y.; Qian, W. Small Object Detection Based on Spatial-Frequency Separated Cross-Attention Swin Transformer. LASER & OPTOELECTRONICS PROGRESS 2025, 62.
- Chen and Ye (2024) Chen, T.; Ye, Z. FreqODEs: Frequency Neural ODE Networks for Infrared Small Target Detection. IEEE TRANSACTIONS ON GEOSCIENCE AND REMOTE SENSING 2024, 62.
- Luo et al. (2024) Luo, Z.; Wang, Y.; Chen, L.; Yang, W. Frequency Spectrum Features Modeling for Real-Time Tiny Object Detection in Remote Sensing Image. IEEE GEOSCIENCE AND REMOTE SENSING LETTERS 2024, 21.
- Li et al. (2024) Li, H.; Yang, J.; Xu, Y.; Wang, R. Mitigate Target-Level Insensitivity of Infrared Small Target Detection via Posterior Distribution Modeling. IEEE JOURNAL OF SELECTED TOPICS IN APPLIED EARTH OBSERVATIONS AND REMOTE SENSING 2024, 17, 13188–13201.
- Liu et al. (2018) Liu, G.; Liu, S.; Wang, F.; Ma, J. Infrared Aerial Small Target Detection with NSCT and Two-Dimensional Property Histogram. INTERNATIONAL JOURNAL OF PATTERN RECOGNITION AND ARTIFICIAL INTELLIGENCE 2018, 32.
- Guo et al. (2025) Guo, H.; Wu, Q.; Wang, Y. AUHF-DETR: A Lightweight Transformer with Spatial Attention and Wavelet Convolution for Embedded UAV Small Object Detection. REMOTE SENSING 2025, 17.
@@
- Zhang and Zhang (2024) Zhang, H.; Zhang, S. Focaler-IoU: More Focused Intersection over Union Loss, 2024, [arXiv:cs/2401.10525].
- Tong et al. (2023) Tong, Z.; Chen, Y.; Xu, Z.; Yu, R. Wise-IoU: Bounding Box Regression Loss with Dynamic Focusing Mechanism, 2023, [arXiv:cs/2301.10051].
- Du et al. (2019) Du, D.; Zhu, P.; Wen, L.; Bian, X.; Ling, H.; Hu, Q.; Peng, T.; Zheng, J.; Wang, X.; Zhang, Y.; et al. VisDrone-DET2019: The Vision Meets Drone Object Detection in Image Challenge Results. In Proceedings of the 2019 IEEE/CVF INTERNATIONAL CONFERENCE ON COMPUTER VISION WORKSHOPS (ICCVW), Seoul, SOUTH KOREA, 2019; IEEE International Conference on Computer Vision Workshops, pp. 213–226.